OpenStudy (anonymous):

Shyla was buying gifts for each of her 8 cousins. She bought everyone either a ring for $6 or a toy for $7. If she spent a total of $53,how many of each did she buy

OpenStudy (anonymous):

She bought 3 ring and 7 toys, because; 7*5=35 3*6=18 18+35=53 5+3=8 Tah-dah! :)